On January 1, 2021, Gundy Enterprises purchases an office building for $173,000, paying $43,000 down and borrowing the remaining $130,000, signing a 9%, 10-year mortgage. Installment payments of $1,646.79 are due at the end of each month, with the first payment due on January 31, 2021.

Can anyone help me figure out how to get this answer?Required information [The following information applies to the questions displayed below.] On January 1, 2021, Gundy Enterpri
Answer #1

Answer

  • Requirement 2

Date

Cash Paid

Interest Expense

Decrease in Carrying Value

Carrying Value

01-Jan-21

$130,000.00

31-Jan-21

$1,646.79

$975.00

$671.79

$129,328.21

28-Feb-21

$1,646.79

$969.96

$676.83

$128,651.38

--Working

Date

Cash Paid

Interest Expense

Decrease in Carrying Value

Carrying Value

01-Jan-21

130000

31-Jan-21

1646.79

=130000*9%*1/12

=1646.79-975

=130000-671.79

28-Feb-21

1646.79

=129328.21*9%*1/12

=1646.79-969.96

=129328.21-676.83